Training Wheel prints custom training material for corporations. The business was started January​ 1, 2017. The company uses a​ normal-costing system. It has two​ direct-cost pools, materials and​ labor, and one​ indirect-cost pool, overhead. Overhead is charged to printing jobs on the basis of direct labor cost. The following information is available for 2017.

Budgeted direct labor costs $225,000

Budgeted overhead costs $247,500

Costs of actual material used $138,000

Actual direct labor costs $216,000

Actual overhead costs $239,600

There were two jobs in process on December​ 31, 2017​:

Job 11 and Job 12. Costs added to each job as of December 31 are as​ follows:

Direct materials

Direct labor

Job 11

$4,810

$5,800

Job 12

$5,630

$7,400

Training Wheel has no​ finished-goods inventories because all printing jobs are transferred to cost of goods sold when completed.

1.

Compute the overhead allocation rate.

2.

Calculate the balance in ending work in process and cost of goods sold before any adjustments for​ under- or overallocated overhead.

3.

Calculate​ under- or overallocated overhead.

4.

Calculate the ending balances in work in process and cost of goods sold if the​ under- or overallocated overhead amount is as​ follows:

a.

Written off to cost of goods sold

b.

Prorated using the overhead allocated 2017
​(before proration) in the ending balances of cost of goods sold and​ work-in-process control accounts

5.

Which of the methods in requirement 4 would you​ choose? Explain.
Answer #1

Pac Answers ③ overhead allocation rate: Let, - BoR = Boco BDLC where, BOL - $247,500 BDLE = $225,000 Now, substitute values iWIP, -Dm +DL ADC OLEDL BOR WIP - DMA DLHOLXDOR - Where, DM = $4,810 + $5, 630 = $10440 DL = $5,800 + $7,400 = $13200 Now, subWow, substitute values in an me = 216,000 + 138,000+2u7,500 me = 601500 The cost of goods sold before adjustment cogs, COGS,u@ The Ending balance or the cost dr goods should @ be Ending labour before adjustment of 560,976-7 560,976 - 7900 = 553076 (2017 year | Budgeted proportion Account ble End balance overhead / Proportion dr before Cost I Loverhead cost afte propractio→ In comparison to cost do goods sold, they ☺ method affects all adjusted accounts that are adjusted to show the changes cons